A bullet of velocity 100m/s can penetrate 2 wooden blocks?If the velocity is doubled ,then how many blocks can it penetrate?

Grade:8

2 Answers

Aditya Gupta
4 years ago

3rd equation of motion can be be applied: v^2 = u^2 + 2as, where acceleration is -ve. Hence v^2 = u^2 – 2as.

v, final velocity = 0. Therefore, u^2 = 2as.

i.e. s is proportional to u^2. → s1÷ s2 = u1^2 ÷ u2^2

Therefore, s2 = (u2^2 ÷ u1^2) x s1 = (200÷100)^2 x 2 (planks) = 8 planks

Answer is 8.

kindly approve :)

Khimraj
4 years ago

When speed is doubled, the kinetic energy is quadruple.

But…

In practice bullets never stay in the same shape when penetrating hard materials, like wood. They deform which alters the path and increases friction…

So theoretically I'd say 8 planks, practically I wouldn't know.